# Domain Specific Pruning of Large Mixture-of-Experts Models with Few-shot Demonstrations

### v1

@@ 4.3 Expert-Level Token Contribution Estimation
When calculating the statistical metrics for evaluating the importance of experts, prior work often directly averages the scores across all tokens [6, 5]. However, in practice, the influence of routed experts’ outputs on the residual stream varies across tokens. As illustrated in Figure 5 (Right), the similarity between representations before and after incorporating the outputs of routed experts differs significantly across tokens for DeepSeek-R1. For some tokens at specific layers, skipping the expert module results in minimal changes to the hidden states, with over 99% similarity between the input and output representations (e.g., the 15th layer of the first tokens). In contrast, for certain tokens and layers (e.g., the 12th layer of the first tokens), the hidden states show substantial differences after expert routing. Intuitively, for the latter type of tokens, modifying their routed experts could introduce a considerable distribution drift in representations [15]. This observation further suggests that such tokens should be given greater weight when assessing the importance of experts.
Inspired by these observations, we propose a similarity-based token importance assessment method. Given the representations before and after the routed expert modules $\bm{h}_{t}^{l}$ and $\tilde{\bm{h}}_{t}^{l}$, we compute the cosine similarity between these representations. The token importance score $s_{t}^{l}$ is then defined as one minus this similarity, capturing the extent of change induced by the routed expert module:
@@ 5.1 Experimental Settings
Evaluated Benchmark. To systematically assess the effectiveness of our proposed method, we conduct experiments across six benchmark datasets: AIME-2024, AIME-2025, HMMT-Feb 2025, LiveCodeBench [14], GPQA-Diamond [13], and AgentBench-OS [16]. These benchmarks encompass four fundamental domains and tasks of LLMs: math, coding, science, and agent-based task execution. We set the maximum context length to 32K, the temperature to 0.6, and the top-p sampling value to 0.95. To ensure statistical reliability, each benchmark is evaluated independently five times, and we report the average performances of pass@1.
Pruning Settings. To demonstrate the domain specialization among DeepSeek-R1’s experts, we perform domain-specific pruning and utilize distinct calibration sets for different domains. For each domain, we randomly sample 25 instances and construct a calibration set by concatenating their inputs with the corresponding outputs generated by DeepSeek-R1. We then evaluate the experts’ scores on the calibration data with our method and select the top 64 and 128 experts with the highest scores at each layer, respectively. We also average the normalized expert scores on different domains to evaluate the mixed-domain pruning performances. Details regarding the candidate sets for each domain are provided in Appendix A.
Baselines. In our experiments, we employ three baseline methods for comparison: (1) Random: A fixed number of experts are randomly selected, while the remaining experts are pruned. (2) Frequency: The Top-$M$ experts with the highest activation frequencies are retained. (3) Gating Score: The Top-$M$ experts with the highest gating scores are retained. The definitions of frequency and gating scores metrics are provided in Equation 3. We do not include perturbation-based methods in our comparison [7, 6], as they require evaluating changes in representations before and after pruning, which is computationally prohibitive for DeepSeek-R1 with 256 experts (the detailed analysis is shown in Appendix B).
@@ 5.2 Main Results
Table 2: Comparison of performances of different expert pruning methods. HMMT denotes HMMT-Feb 2025, GPQA denotes GPQA-Diamond, and A-OS denotes AgentBench-OS.
[…]
Full - 256 78.00 65.33 45.33 63.32 70.91 40.51 60.57 -
Random $\times$ 64 0.00 0.00 0.00 0.00 26.09 0.00 4.35 7.18
Frequency $\times$ 64 0.00 0.00 0.00 0.00 17.68 2.78 3.41 5.63
Gating Score $\times$ 64 2.67 1.33 2.67 0.00 20.20 0.69 4.59 7.58
EASY-EP $\times$ 64 72.67 55.33 36.00 42.51 67.47 27.26 50.20 82.88
Random $\times$ 128 8.33 6.67 3.33 20.96 41.41 7.64 14.72 24.30
Frequency $\times$ 128 19.33 13.33 7.33 16.17 60.01 29.16 24.22 39.99
Gating Score $\times$ 128 69.33 58.67 37.33 46.71 62.63 31.94 51.10 84.37
EASY-EP $\times$ 128 80.67 66.67 48.33 61.11 70.12 37.92 60.80 100.38
Frequency 128 18.00 13.33 8.00 9.58 30.30 13.19 15.40 25.42
Gating Score 128 18.67 8.67 8.67 53.29 37.37 25.00 25.28 41.73
EASY-EP 128 80.67 62.00 39.33 58.38 70.51 34.72 57.60 95.10
Table 2 shows the performance comparison of our method with baselines on different pruning settings. Firstly, our method outperforms other baselines and even surpasses the full DeepSeek-R1 model with only half of the experts on certain datasets for domain-specific pruning. Across all benchmarks and pruning settings, our approach consistently achieves significantly better results than all baselines. Moreover, under domain-specific pruning settings, our method maintains performance comparable to full DeepSeek-R1 and even exceeds it on math-related benchmarks. We hypothesize that removing irrelevant experts allows the model to rely more on domain-specific knowledge, leading to improved performance.
Secondly, our method can maintain performance even with a large compression ratio, as shown in Figure 1. When the compression ratio is large (e.g., $75\%$), most pruning methods can hardly correctly complete the query at all, manifested as 00 scores on many datasets. Additionally, they even fail to follow instructions correctly and produce nonsensical speech, e.g., failing to generate option letters for GPQA. On the contrary, our method can preserve $82.88\%$ averaged performances across all datasets and even $95.15\%$ performances on GPQA.
@@
Finally, our method can also preserve performances under mixed-domain pruning. By leveraging mixed-domain pruning, our approach effectively preserves performance across multiple target domains, retaining $95.09\%$ of the original performance, which even surpasses domain-specific pruning methods via gating score metric. In contrast, other expert pruning techniques struggle to maintain a balanced performance across different domains. This result highlights that the overlapped experts identified by our method, which are associated with general reasoning abilities, can effectively contribute to a wide range of downstream domains.
@@ 5.3 Detailed Analysis
In this section, we conduct further ablation studies and detailed analyses to investigate the effectiveness of our approach and few-shot expert identification phenomena of DeepSeek-R1.
@@ 5.3.1 Ablation Study
Ours $g_{i,t}^{l}\lVert\bm{e}_{i,t}^{l}\rVert\cdot s^{l}_{t}$ 64 72.67 55.33 36.00 42.51 67.47 27.26
@@ 5.3.2 Effect of Pruning Data
We also explore the effect of pruning data. Instead of the input and output of full DeepSeek-R1, we consider full different data: (1) Inp: only input data; (2) R1: only generation data by DeepSeek-R1; (3) Inp+Ans: input data and golden answer; and (4) PT: pre-training data within the same domain. The experiment results are displayed in Table 4. Compared to using the concatenation of input and DeepSeek-R1’s output, the model’s performance under other settings suffered degradation in both math and coding tasks. This indicates that even within the same domain, distinct differences of expert distributions exist between the input data, golden answer, the model’s reasoning processes, and the pretraining data.
5.3.3 Generalization Capacity
In addition to evaluating performance on tasks within the same domain as the pruning data, we also assess tasks from unrelated domains (e.g., pruning DeepSeek-R1 with AIME2023 data while evaluating on LiveCodeBench). As presented in Table 5, out-of-domain performance declines compared to in-domain results but still retains a notable level of generalization capability. We hypothesize that although some experts associated with specific domain knowledge were pruned, experts related to core reasoning abilities were preserved. Specifically, the model demonstrates better generalization performance between math and science domains, which may be attributed to the inherent conceptual similarities between these fields.
Table 4: Comparison of performances with different pruning data.
@@ 5.3.3 Generalization Capacity
Table 5: Results of generalization capacities of pruned models. Domain denotes the domain of pruning data.
Domain AIME24 LiveCodebench GPQA
Math 80.67 46.11 46.91
Coding 38.00 61.11 39.90
Science 64.64 53.59 70.12
Figure 6: Left: Comparison of performances with different number of shots for pruning. Right: Total Throughput across different numbers of experts.
@@ 5.3.4 Effect of Number of Pruning Demonstrations
In Section 3.2, we demonstrated that domain-relevant experts can be effectively identified using only a few demonstrations. Here, we further investigate the impact of the number of demonstrations on final performance. To do so, we sample varying numbers of demonstrations from the same distribution and prune half of the experts in each layer. The performance variations on AIME24 and LiveCodeBench are presented in Figure 6 (Left). When we utilize only a single sample for pruning, the selected experts are often influenced by the characteristics of that individual sample, thus resulting in lower performance. As we further increase the number of demonstrations, the performance rapidly rises, achieving comparable performances with the full model.
5.4 Analysis of Throughout
To evaluate the throughput of pruned models with different numbers of experts, we use the SGLang [17] package and measure performance under a maximum request concurrency of 32. We consider four settings: (1) 1K+1K: 1K input length and 1K output length; (2) 1K+4K: 1K input length and 4K output length; (3) 4K+1K: 4K input length and 1K output length; and (4) 4K+4K: 4K input length and 4K output length. For configurations with no more than 192 experts, we use a single 8$\times$×H800 GPU and double the measured throughput. For configurations with more than 192 experts, two 8$\times$×H800 GPUs are used. Figure 1 shows the scaling throughput for the 1K+1K setting, while Figure 6 (Right) presents results for the other settings.
We observe that reducing the number of experts significantly improves throughput, particularly when the model can be deployed on a single node. Compared to the full DeepSeek-R1 model, configurations with 128 and 64 experts achieve 2.99$\times$ and 4.33$\times$ higher throughput, respectively, under the 1K+1K setting. Under the 4K+4K setting, they reach 2.51$\times$ and 2.97$\times$ improvements. Furthermore, in long output scenarios with an equivalent total sequence length, the 128-expert configuration achieves a significantly higher acceleration ratio (2.91$\times$ under the 1K+4K setting, compared to 2.52$\times$ under the 4K+1K setting).
@@ 6 Related Work
6.1 Sparse MoE Architectures
Sparse MoE architectures aim to improve computational efficiency by activating only a small subset of experts for each input [18, 19, 4]. These methods focus on designing inherently efficient MoE architectures, leveraging conditional computation to scale up model capacity without linearly increasing inference cost. Driven by scaling laws, the parameter size of dense models has grown dramatically, leading to the emergence of even larger MoE models [20, 2, 21]. To further enhance the efficiency of MoE, DeepSpeed-MoE [22] proposes a pyramid-like architecture where the lower layers contain more experts. Similarly, DeepSeek-MoE [10] adopts a fine-grained expert design that reduces the parameter size per expert while increasing the total number of experts. In addition, another line of research focuses on transforming dense models into MoE architectures to reduce computational costs [23, 24]. By introducing sparsity into existing dense architectures, these methods aim to retain the performance of large models while benefiting from the conditional computation property of MoE. These studies primarily concentrate on architectural design. In contrast, another line of research aims to compress already-trained MoE models using post-hoc methods such as pruning and quantization, which we discuss in the next section.
6.2 MoE Compression Methods
While sparse MoE architectures improve computational efficiency by activating only a small subset of experts, they still incur significant memory overhead due to the need to store all expert parameters [25], posing a different set of challenges from existing dense LLM compression [26, 27]. The prior study [28] has observed that many experts contain redundant information, motivating efforts to further reduce memory usage by eliminating unnecessary experts. A common approach is to apply sparsification techniques to retain only the most important experts, based on carefully curated importance metrics. These metrics include the frequency [5] with which experts are selected during routing (i.e., “heavy hitter counts”), as well as analyses of expert gating scores and activation patterns [29, 5]. In addition to selection-based strategies, parameter sharing among experts can also be employed to alleviate the overall memory overhead [28, 30]. Some studies further assess expert importance by evaluating their impact on the overall model output, aiming to search for the optimal combination of experts [6, 7]. However, such methods involve exploring a large combinatorial search space, which leads to significant computational overhead and makes them impractical for large-scale MoE models. Additionally, another line of research leverages expert representations and router logits to merge similar experts into a single expert, thereby reducing the total number of experts [31, 32]. However, these methods introduce additional parameters and incur significant memory costs, particularly when performing multiple task-specific pruning. Different from them, our approach enables efficient pruning with just a single forward pass and eliminates the need for extra memory to store new models.
@@ 7 Conclusion
In this work, we investigated the domain specialization of experts in DeepSeek-R1. Our observations indicate that domain-specific experts play a crucial role in their respective domains and can be effectively identified with a few demonstrations. Building on these insights, we proposed a pruning strategy that leverages demonstrations from tasks within the same domain. Specifically, we introduced EASY-EP, a simple yet effective pruning method that combines output-aware expert importance assessment with expert-level token contribution estimation. Experimental results showed that our approach maintained comparable performance while utilizing only half of the experts in domain-specific settings and retains approximately $95\%$ of the original performance in mixed-domain pruning. We believe that our method can facilitate the deployment of large MoE models, particularly for efficiently handling a high volume of samples within the same domain.
Appendix A Experiment Details

### v2

@@ 4.3 Expert-Level Token Contribution Estimation
When calculating the statistical metrics for evaluating the importance of experts, prior work often directly averages the scores across all tokens (6; 5). However, in practice, the influence of routed experts’ outputs on the residual stream varies significantly across tokens (as shown in Appendix D.2). Intuitively, when dealing with tokens exhibiting low similarity before and after the MoE module, adjusting their routed experts will induce a substantial distributional shift in their representations. In contrast, for tokens with high similarity, such adjustments will lead to only minimal drift in their representational distributions (20). Inspired by these, we propose a similarity-based token importance assessment method which give greater weights for the former tokens. Given the representations before and after the routed expert modules $\bm{h}_{t}^{l}$ and $\tilde{\bm{h}}_{t}^{l}$, we compute the cosine similarity between these representations. The token importance score $s_{t}^{l}$ is then defined as one minus this similarity, capturing the extent of change induced by the routed expert module:
@@ 5.1 Experimental Settings
Evaluation Benchmarks. To systematically assess the effectiveness of our proposed method, we conduct experiments across eight benchmark datasets: AIME-2024, AIME-2025, HMMT-Feb 2025, LiveCodeBench (19), GPQA-Diamond (18), USMLE (21), FinanceIQ (22), and AgentBench-OS (23). These benchmarks encompass six fundamental domains and tasks of LLMs: math, coding, science, medicine, finance, and agent-based task execution.
Experiment Settings. We select DeepSeek-R1 (1) and DeepSeek-V3-0324 (15) as our evaluated models and consider domain-specific and mixed-domain pruning settings. For each domain, we randomly sample 25 instances and construct a calibration set by concatenating their inputs with the target model’s outputs. We then evaluate the expert scores on the calibration data and select the top 64 and 128 experts with the highest scores at each layer, respectively. We also average the normalized expert scores on different domains to evaluate the mixed-domain pruning performances. Details regarding the candidate sets and evaluation settings are provided in Appendix E.
Baselines. In our experiments, we employ three expert pruning methods and one expert merging method for comparison. For expert pruning, we employ different methods to assess the expert scores, including random, frequency, and gating scores (as discussed in Equation 2), and only keep Top-$M$ experts with the highest expert scores  [^2] We do not include perturbation-based pruning methods in our comparison (6; 7), which is computationally prohibitive for MoE models with 256 experts (the detailed analysis is shown in Appendix F).. For expert merging, we select M-SMoE (24), which first employs neuron permutation alignment to mitigate neuron misalignment of experts and then merges experts into dominant ones with similarity of router logits.
@@ 5.2 Main Results
Table 2: Comparison of the performances of different expert pruning methods. HMMT denotes HMMT-Feb 2025, GPQA denotes GPQA-Diamond, A-OS denotes AgentBench-OS, and FinIQ denotes FinanceIQ.
Model Method Mix #E AIME-24 AIME-25 FMMT LiveCode GPQA USMLE FinIQ A-OS Avg
DeepSeek -R1 Full - 256 77.08 66.67 44.38 63.32 70.91 92.66 82.1 40.51 67.20
Random $\times$ 64 0.00 0.00 0.00 0.00 26.09 0.00 0.00 0.00 3.26
Frequency $\times$ 64 0.00 0.00 0.00 0.00 17.68 0.00 0.00 2.78 2.58
Gating Score $\times$ 64 2.67 1.33 2.67 14.97 46.83 0.86 0.00 0.69 8.75
M-SMoE $\times$ 64 0.00 0.00 0.00 0.00 12.12 0.00 0.00 0.00 1.52
EASY-EP $\times$ 64 72.81 55.10 38.02 42.51 67.47 26.63 33.90 27.26 45.22
Random $\times$ 128 8.33 6.67 3.33 20.96 34.95 57.66 0.00 7.64 17.44
Frequency $\times$ 128 19.33 13.33 7.33 36.08 59.60 61.51 26.40 29.16 31.59
Gating Score $\times$ 128 70.10 55.52 36.15 47.60 63.78 80.36 66.50 31.94 56.49
M-SMoE $\times$ 128 5.33 6.00 3.33 25.75 24.75 52.63 39.60 19.44 22.10
EASY-EP $\times$ 128 79.17 68.33 45.31 61.11 70.12 91.67 78.80 37.92 66.55
Frequency 128 21.33 10.00 6.00 7.49 41.45 78.55 62.14 11.81 29.85
Gating Score 128 29.33 21.33 18.00 22.75 41.69 62.06 27.29 30.56 31.67
M-SMoE 128 6.67 2.00 4.67 4.19 32.32 72.00 19.10 6.25 18.40
EASY-EP 128 75.94 61.98 42.50 57.63 70.36 91.20 57.95 34.17 61.47
DeepSeek -V3-0324 Full - 256 55.73 47.71 28.75 48.50 66.87 87.51 64.22 33.33 54.08
Random $\times$ 64 0.00 0.00 0.00 0.00 26.87 0.39 0.00 0.69 3.49
Frequency $\times$ 64 31.35 34.06 15.73 1.95 45.25 40.13 61.96 22.74 31.65
Gating Score $\times$ 64 43.96 25.10 23.12 14.97 51.52 78.68 64.20 0.00 37.69
M-SMoE $\times$ 64 16.67 13.33 3.33 1.20 22.22 12.18 47.00 21.52 17.18
EASY-EP $\times$ 64 53.12 41.56 28.85 27.99 57.35 84.57 72.50 27.55 49.19
Random $\times$ 128 1.33 0.67 0.00 11.38 34.95 53.5 53.66 18.75 21.78
Frequency $\times$ 128 55.73 42.60 30.10 36.08 63.54 84.29 66.84 31.71 51.36
Gating Score $\times$ 128 55.42 45.10 30.94 47.60 63.78 84.62 67.76 35.42 53.83
M-SMoE $\times$ 128 48.00 38.67 28.67 30.53 55.82 86.72 66.60 33.33 48.54
EASY-EP $\times$ 128 55.21 46.88 31.56 46.71 65.25 86.72 63.58 37.08 54.12
Frequency 128 51.35 37.60 24.27 17.07 55.90 83.47 66.80 36.25 46.59
Gating Score 128 53.75 40.10 27.19 28.74 58.88 83.86 67.74 34.58 49.36
M-SMoE 128 43.33 30.00 20.00 7.19 52.53 82.33 62.20 29.17 40.84
EASY-EP 128 57.81 46.56 33.33 40.72 64.95 85.00 72.26 38.74 54.92
Table 2 showcases our method’s performance against baselines under various pruning configurations. First, our approach consistently outperforms all baseline methods across diverse benchmarks and pruning settings. Notably, in domain-specific pruning, our method matches and even surpasses full model performance on certain benchmarks with only half the experts. This may be attributed to the effective removal of irrelevant experts, enhancing the model’s ability to utilize domain-specific knowledge. Furthermore, our method demonstrates strong resilience to high compression ratios (e.g., $75\%$), where most methods experience significant performance degradation, particularly on DeepSeek-R1. In contrast, our technique preserves substantial model capabilities, highlighting its effectiveness in identifying critical experts for specific tasks.
Second, non-reasoning models exhibit greater robustness compared to their reasoning-oriented counterparts. DeepSeek-V3-0324, for instance, retained more performance across most pruning methods than DeepSeek-R1. Under domain-specific pruning, DeepSeek-V3-0324’s performance on datasets like AgentBench-OS even improved notably after pruning some experts. However, this phenomenon is not observed with DeepSeek-R1. We hypothesize that while domain capabilities might be preserved in pruned reasoning models, their long-term generation abilities are compromised.
@@
Finally, our method also excels in preserving performance under mixed-domain pruning. It retains over $90\%$ of the original performance, surpassing domain-specific compression with other methods. Conversely, other expert pruning techniques struggle to maintain balanced performance across different domains. This underscores that the overlapped experts identified by our approach, which are linked to general reasoning abilities, effectively contribute to a broad array of downstream domains.
@@ 5.3 Detailed Analysis
In this section, we conduct further ablation studies and detailed analyses to investigate the effectiveness of our approach and the few-shot expert localization phenomena of large MoE models. For more analysis experiments, we present them in Appendix G.
@@ 5.3.1 Ablation Study
Ours $g_{i,t}^{l}\lVert\bm{e}_{i,t}^{l}\rVert\cdot s^{l}_{t}$ 64 72.81 55.33 36.00 42.51 67.47 27.26
@@
5.3.2 Generalization Capacity
Beyond same-domain task evaluations, we assessed the generalization capacities on unrelated domains after domain-specific pruning (e.g., pruning DeepSeek-R1 with AIME2023 and evaluating on LiveCodeBench). As Table 4 shows, the model demonstrates a certain generalization ability, especially in similar domains (e.g., math/science, code/OS-agent, science/medicine). We hypothesize that while some domain-specific experts are pruned, core reasoning experts are preserved. However, the out-of-domain performances are typically poorer than mixed-domain pruning. Thus, we suggest using mixed-domain pruning when facing multiple downstream tasks.
Table 4: Results of generalization capacities of pruned models. Domain denotes the domain of pruning data. Bold denotes in-domain performances.
Domain AIME24 LiveCodeBench GPQA Agent-OS USMLE FinIQ
Math 79.17 46.11 46.91 3.47 46.43 58.20
Coding 38.00 61.11 39.90 15.97 41.79 53.00
Science 64.64 53.59 70.12 4.17 75.88 57.50
@@ 5.3.3 Effect of Pruning Data
We also study the impact of pruning data. Instead of using the full model’s input and output, we examine five types of data: (1) Inp: just input context data; (2) Out: only the model’s generated data; (3) Inp+Ans: input context data and the correct generated answer; (4) PT: pre-training data from the same domain; and (5) CC: data from CommonCrawl. Table 5 and 6 show the experimental results for DeekSeek-R1 and DeepSeek-V3-0324. Compared to using the combination of input and model output, performance decreased under other settings for both math and coding tasks. This suggests that, even within the same domain, there are notable differences in expert distributions among data types. Additionally, employing CommonCrawl data causes significant performance declines, demonstrating that pruning a task-irrelevant model with pre-training data is not suitable for current models.
Table 5: Comparison of performances of DeepSeek-R1 with different data.
@@
CC 0.00 0.00 28.92
Table 6: Comparison of performances of DeepSeek-V3-0324 with different data.
Data AIME24 LiveCodebench GPQA
Inp+Out 55.73 48.50 66.87
Inp+Ans 54.00 43.11 61.62
Inp 51.33 14.97 63.64
Out 54.67 47.90 63.13
PT 13.33 37.72 56.57
CC 0.00 0.00 29.80
Figure 4: Comparison of performance with different numbers of shots for pruning.
@@ 5.3.4 Effect of Number of Pruning Demonstrations
In Section 3.2, we observe that domain-relevant experts can be identified with only a few demonstrations. Here, we further investigate the impact of the number of demonstrations on final performance. To do so, we sample varying numbers of demonstrations from the same distribution and prune half of the experts in each layer. The performance variations on AIME24 and LiveCodeBench are presented in Figure 4. When we utilize only a single sample for pruning, the selected experts are often influenced by the characteristics of that individual sample, thus resulting in lower performance. As we further increase the number of demonstrations, the performance rapidly rises, achieving comparable performances with the full model.
5.3.5 Analysis of Throughput
To evaluate the throughput of pruned models with different numbers of experts, we use the SGLang (25) package and measure performance under a maximum request concurrency of 32. We evaluate the settings with 1K input and 1K output length. For configurations with more than 192 experts, two 8$\times$×H800 GPUs are used. Figure 1 shows the scaling throughput for this setting. We observe that reducing the number of experts significantly improves throughput, particularly when the model can be deployed on a single node. Compared to the full DeepSeek-R1 model, configurations with 128 and 64 experts achieve 2.99$\times$ and 4.33$\times$ throughput, respectively. Compared to the full model, the pruned model can be deployed on a single node, thereby avoiding inter-node communication overhead. Moreover, using fewer experts further reduces communication between GPUs within the node, improving computational efficiency. We provide more experiments in Appendix G.3.
@@ 6 Related Work
MoE architectures improve computational efficiency by activating only a small subset of experts per input (26; 27; 4). However, the increased number of parameters introduced by these architectures leads to substantial memory overhead. To address this, existing efforts broadly fall into two categories. The first line of work focuses on architectural optimization to reduce computation or parameter size. Representative methods include pyramid-shaped expert allocation (28), fine-grained expert design with smaller per-expert modules (14), and the transformation of dense models into sparse MoE variants (29; 30). While effective, these approaches typically require modifying model architecture and retraining from scratch, which limits their applicability to already deployed or pretrained models. The second line of research approaches the problem from a memory efficiency perspective by applying post-hoc compression techniques, primarily pruning and quantization (31; 32). These methods typically estimate expert importance based on routing frequency (5), gating scores (33), or direct measurement of their contribution to model outputs (6; 7). Some approaches further reduce redundancy by merging similar experts (8; 9), though they may meet the problem of neuron misalignment and additional memory costs (10; 11). In contrast, our method enables efficient pruning with a single forward pass and avoids storing additional model variants during compression.
@@ 7 Conclusion
In this work, we investigated the domain specialization of experts in large MoE models. Our observations indicate that domain-specific experts play a crucial role in their respective domains and can be effectively identified with a few demonstrations. Building on these insights, we proposed a pruning strategy that leverages demonstrations from tasks within the same domain. Specifically, we introduced EASY-EP, a simple yet effective pruning method that combines output-aware expert importance assessment with expert-level token contribution estimation. Experimental results showed that our approach maintained comparable performance while utilizing only half of the experts in domain-specific settings and retained over $90\%$ of the original performance in mixed-domain pruning. We believe that our method can facilitate the deployment of large MoE models, particularly for efficiently handling a high volume of samples within the same domain.
Appendix A Limitation
Our work investigates the phenomenon of few-shot expert localization in large MoE models and proposes a simple yet effective method for domain-specific expert pruning. We leave the investigation of training to further enhance the pruned model’s performance, particularly in balancing in-domain and out-of-domain capabilities, as future work, given our current focus on evaluating pruning effectiveness under realistic resource constraints. We observed differing levels of robustness to expert pruning between reasoning-oriented (DeepSeek-R1) and non-reasoning models (DeepSeek-V3-0324). While this trend is noteworthy, further validation on a broader range of architectures is needed to strengthen the generality of this observation.
Appendix B Expert Overlap Across Domains
To further analyze the overlap of top experts across different domains, we select AIME-2023, LiveCodeBench-V3, and GPQA as calibration sets and select top experts with gating scores. Subsequently, we compute the overlap ratio of Top-$M$ experts ($M\in\{1,2,4,8,16,32,64,128\}$) and expert overlap across different layers, which are shown in Figure 5. We can observe that the experts with the highest gating scores differ significantly across different domains, and as the number of top experts increases, the overlap ratio increases. Additionally, in the lower layers, there is a higher overlap among experts with significant gating scores across different datasets, but this overlap becomes relatively lower in the middle and deeper layers. This indicates that experts in the lower layers tend to focus on general capacities, while as the network depth increases, they gradually specialize in handling knowledge from distinct domains.
Figure 5: Left: Overlap ratios of Top-$M$ experts with different $M$. Middle and Right: Overlap experts in top-16 and 128 experts on different layers.
Appendix C Empirical Study with EASY-EP
To further illustrate the phenomenon of few-shot expert localization, we analyze it with the expert scores (as discussed in Equation 3) from our proposed EASY-EP method in place of the gating scores.
C.1 Expert Specialization Across Domains
Distinct Expert Distribution Across Domains. We compute the overlap ratios of the top experts identified by our method and visualize them in Figure 6. Similar to previous observations, large differences in expert distribution still exist across different domains. However, the expert overlap between domains is larger than that of experts identified by gating scores. This may be because our method identifies experts not only frequently activated within a domain but also those who contribute more than other experts in residual connections. The latter may be more similar across different datasets, thus leading to higher overlap.
Figure 6: Overlap ratios of experts with top expert scores computed with our methods on different datasets.
Impact of Removing Domain-Specific Experts. Similarly, we remove experts whose score, as calculated by our method, falls within the top-128 for a single domain but not for any other. As shown in Table 7, removing these in-domain experts leads to larger performance degradation than out-of-domain experts. However, compared with pruning with gating scores, the performance degradation is less severe. We speculate this is due to the higher overlap among the top experts identified by our method, resulting in a smaller pruning ratio.
Table 7: Results of removing domain-specific experts with EASY-EP.
Domain AIME24 GPQA LiveCodeBench
Full 77.08 70.91 63.32
Math 72.67 (-3.41) 73.17 (+2.26) 64.22 (+0.90)
Code 76.67 (-0.41) 72.93 (+2.02) 62.27 (-1.05)
Science 76.00 (-1.08) 58.23 (-12.68) 64.97 (+1.65)
C.2 Expert Locality Within One Domain
Figure 7: Left: Overlap ratio of top-128 experts identified by EASY-EP with different numbers of demonstrations. Right: Overlap ratio of top-128 experts identified by EASY-EP pruned with different math datasets.
Effect of Calibration Set Size. We present the overlap ratios of top-128 experts selected via our method with different numbers of demonstrations. As shown in Figure 7 (Left), as the number of demonstrations increases, the identified domain-specific experts gradually become stable. For 25-shot pruning, it can achieve the overlap ratio of $93\%$ with 10-shot pruning, further demonstrating the feasibility of few-shot domain-specific expert pruning.
Consistency of Expert Activation Across Datasets. Figure 7 (Right) illustrates that the top-128 experts identified by our method exhibit consistent overlap ratios across different datasets within the math domain. Our method not only presents expert consistency over different datasets similar to the gating score metric, but also yields significantly higher overlap ratios. We hypothesize that our approach is more adept at identifying experts of true domain-wide significance, as opposed to those who are only important in a single dataset due to the dataset-specific differences.
Appendix D Empirical Analysis of Components in EASY-EP
D.1 Output-Aware Expert Importance Assessment
To analyze the relationship between whether a large gating score $g_{i,t}^{l}$ ensures a large product of gating scores and L2 norm of expert output $g_{i,t}^{l}\lVert\bm{e}_{i,t}^{l}\rVert$, we visualize the relationship of top-128 experts selected by the two metrics. As shown in Figure 8 (Left), despite a significant degree of overlap among experts, there still exist some experts who excel exclusively in a single, focused metric. Although the gating scores of some experts are not large, the L2 norms of their outputs are larger than others. This proves the necessity of considering experts’ outputs.
D.2 Expert-Level Token Contribution Estimation
Given one sample selected from AIME-2023, we first obtain representations before and after incorporating the outputs of routed experts in each layer of DeepSeek-R1 and compute the similarities between them, as illustrated in Figure 8 (Right). We can observe that the similarities differ significantly across tokens and layers. For some tokens at specific layers, skipping the expert module results in minimal changes to the hidden states, with over $99\%$ similarity between the input and output representations (e.g., the 15th layer of the first tokens). In contrast, for certain tokens and layers (e.g., the 12th layer of the first tokens), the hidden states show substantial differences after expert routing.
Appendix E Experiment Details